a map of Indiana is being made to a scale of 1/2 in: 20 mi. Indianapolis is 128 miles from fort wayne. how far apart from each other will they appear on this map?
Answer by rfer(16322)

0.5/20=x/128
20x=64
x=64/20
x=3.2 inches